WHAT HAPPENS TO THE AREA WHEN THE LENGTH OF A SIDE OF A SQUARE IS DOUBLED?..
WHAT HAPPENS TO THE AREA WHEN THE LENGTH OF A SIDE IS TRIPLED?..

Answer by stanbon(75887)

WHAT HAPPENS TO THE AREA WHEN THE LENGTH OF A SIDE OF A SQUARE IS DOUBLED?..
Before change: Area = L*W
After change: Area = (2L)*W = 2(L*W)
So the area has doubled.
---------------------------
WHAT HAPPENS TO THE AREA WHEN THE LENGTH OF A SIDE IS TRIPLED?..
Before change : Area = L*W
After change: Area = (3L)*W = 3(L*W)
So the area is 3 times what it was originally
=======================
Cheers,
Stan H.